# Open3DTrack: Towards Open-Vocabulary 3D Multi-Object Tracking

### v1

@@ I INTRODUCTION
In this paper, we introduce the task of open-vocabulary 3D tracking, where trajectories of known and unknown object classes are estimated in the real 3D space by linking their positions across consecutive frames. We formulate the open-vocabulary tracking problem, propose various splits to evaluate the performance across known and unknown classes, and introduce a novel approach that bridges the performance gap between tracking known and unseen classes. Our approach leverages the strengths of 2D open-vocabulary methods and existing 3D tracking frameworks to generalize to unseen categories effectively. Through extensive experiments, we demonstrate that our proposed method achieves strong results through our adaptation strategies, providing a robust solution for open-vocabulary 3D tracking. Fig. 1 shows an example output of our solution for the open-vocabulary 3D tracking task. Our method can detect and track previously unseen object categories in three-dimensional space without relying on predefined labels.
@@ II Related Work
In 3D object tracking, various approaches have been explored, each leveraging different features and methodologies. Traditional approaches often rely on Kalman filters or other probabilistic models, with recent advances shifting towards deep learning-based methods. The tracking-by-detection paradigm remains dominant, particularly in 3D multi-object tracking (MOT). Model-based techniques, such as AB3DMOT [7], extend traditional Kalman filters to 3D for motion estimation, coupling it with the Hungarian method to provide a standard 3D MOT baseline. Another method, EagerMOT [8], combines observations from LiDAR and camera sensors, enhancing the ability to track distant objects and achieve precise trajectory localization. The Immortal Tracker [9] introduces a trajectory prediction mechanism using a Kalman filter to maintain tracks even when objects temporarily disappear from view, significantly reducing identity switches. CenterPoint [3] and CenterTrack [10] combine velocity estimates with point-based detection to track the centers of objects across multiple frames. Poly-MOT [1] introduces a method that adapts a tracking criterion based on object categories, employing category-specific motion models and a two-stage data association strategy. In 3DMOTFormer [2], the authors introduce a geometry-based 3D multi-object tracking framework built on the transformer architecture, using an Edge-Augmented Graph Transformer to perform data association through edge classification. While great advancement has been made in the 3D tracking task, all these methods rely on tracking a known predefined set of object classes.
@@ II-B Open-Vocabulary 3D
Open vocabulary 3D tasks are relatively new and emerging in computer vision, driven by the need to generalize beyond fixed, predefined categories, particularly in dynamic and unstructured environments. Recent advances in 2D open vocabulary tasks [11, 12, 13, 14] have inspired the extension of these concepts to the 3D domain, where the complexity of object shapes, occlusions, and sparse point-cloud data pose unique challenges. Introducing open vocabulary techniques into 3D vision tasks marks a significant shift towards more flexible and scalable models capable of identifying objects across a wide range of categories without requiring exhaustive labeled data. In [15], one of the earlier works, the authors applied transductive zero-shot learning (ZSL) and generalized zero-shot learning (GZSL) to 3D point cloud classification. PointCLIP [16] leverages the Contrastive Vision-Language Pre-training (CLIP) model to achieve open-vocabulary point cloud recognition by aligning CLIP-encoded point clouds with 3D category texts through knowledge transfer from 2D images to 3D. For 3D open-vocabulary tasks, using image features obtained from vision language models and aligning them with point cloud features is common [17], [18]. OpenScene [19] generates dense features for 3D points embedded within the CLIP feature space by using image-per-pixel features from 2D semantic segmentation models [13]. OpenMask3D [20] introduces a zero-shot approach for open-vocabulary 3D instance segmentation by utilizing class-agnostic 3D instance masks and aggregating per-mask features through multi-view fusion of CLIP-based image embeddings. In addition to semantic segmentation, in [21], the authors propose a method that combines a point cloud detector for object localization with cross-modal contrastive learning to connect image, point cloud, and text representations, enabling open-vocabulary 3D detection in indoor scenes. However, open-vocabulary 3D tracking remains largely unexplored, presenting a significant gap in leveraging these emerging techniques for dynamic, real-time object tracking in the physical space.
@@ III-A Problem Formulation
Given a sequence of LiDAR frames $\mathbf{P}=\{P_{t}\}_{t=1}^{T}$, where $P_{t}$ represents the point cloud at time $t$, the goal of 3D multi-object tracking is to detect and track multiple objects $\{O_{i}^{t}\}_{i=1}^{N_{t}}$, where $O_{i}^{t}$ denotes the $i$-th object in frame $t$. Each $O_{i}^{t}$ comprises the object class $c$ and the 3D bounding box $b=[x,y,z,w,l,h,\theta]$ where $x,y,z$ represent the coordinates of the center of the bounding box, and $w,l,h$ are the width, length, and height of the bounding box, and $\theta$ is the heading angle, respectively.
@@ III-B Open-Vocabulary 3D Multi-Object Tracking
While 3DMOTFormer performs well for the closed set, it struggles to track unseen object classes due to its reliance on class-specific information during training. The tracker expects identified 3D boxes to generate edge proposals and then process those through the graph transformer. In a situation where it may receive a new label or an unlabelled object, the model will fail to track it. 3DMOTFormer also uses class-specific statistics within its pipeline, such as maximal class velocities, to compute association edges between tracks and new detections. Such information regarding unseen, unfamiliar objects, that the system may encounter for the first time, is not available in a real-world open-vocabulary setting.
Class Agnostic Tracking
To enable tracking objects irrespective of their class label, we modify the tracker to utilize only object position, dimensions, heading angle and velocity as the initial features. Object detections from a given detector are treated as proposals (class-agnostic) where the class labels and confidence scores are dropped and not utilized.
To compute potential association edges, we replace class-specific distance thresholds with a fixed threshold, achieving a balance between maintaining true edges and avoiding unnecessary associations.
We change the GT track ID assignment by computing 3D IoU between ground truth boxes $A^{train}$ and 3D detections belonging to $\mathcal{C}^{base}$ in a class-agnostic manner.
@@
As discussed earlier, we drop class labels and confidence scores from the detection features to ensure the tracker is independent of any information specific to object classes. As a result, the initial output from the system suffers from a lack of representation of the objectness score of each box. To overcome this issue, we use the confidence scores available for $\mathcal{C}^{base}$ and train a regression head at the end of the tracker. The regression head is composed of a feed-forward network that takes as input processed detection features $h^{(Ld)}_{D,j}$ from the graph transformer and predicts a single confidence score. We employ a straightforward Mean Squared Error loss between predicted and true confidence scores to learn the scores for the detection boxes. The combined loss is $\mathcal{L}=\mathcal{L}_{m}+\lambda_{c}\mathcal{L}_{c}$, where $\mathcal{L}_{c}$ is the loss for confidence score, weighted by $\lambda_{c}$.
@@
While we train the tracker by removing all class-specific dependencies, we require a strong open-vocabulary classification for all object tracks, known and unknown, at test time. We leverage vision-language models to match regions in the images with text prompts, enabling it to detect and classify objects for which it has not been explicitly trained. 2D detections on multi-view images from the vehicle are obtained from an open-vocabulary 2D detector, which is prompted with all possible labels—both $\mathcal{C}^{base}$ and $\mathcal{C}^{novel}$. The system allocates class labels to the 3D proposals by first projecting the 3D boxes onto multi-view images. The labels are then derived from the highest IoU overlap with 2D detections on the image plane. Projections of the 3D detections that do not overlap with a 2D open-vocabulary detection are labeled ’unknown’. We explicitly identify these boxes at the end of the tracking process.
@@
At the output of the tracker, we apply a track smoothing module that addresses the issue of unknown boxes that do not match any 2D detections from the images, as well as improve label consistency throughout tracks to account for inaccuracies in the open-vocabulary detector outputs and errors accrued during projection from 3D to 2D. To calculate the most accurate class for each track, we first compute the weight of an object based on its detected bounding box in an image, with objects at a greater distance assigned a smaller weight. The depth is derived from the size of the bounding box, the object’s vertical position in the image for perspective correction, and the aspect ratio of the bounding box to ensure uniformity for all classes.
$B_{size}=\frac{B_{width}\times B_{height}}{I_{width}\times I_{height}}$
@@
where, $B$ refers to 2D detected box, $I$ refers to image, $y_{center}$ points to the vertical center of the box, $\alpha_{p}$ is the perspective correction factor, while $\beta_{ar}$ is the aspect ratio threshold and
$aspect\_ratio=\frac{B_{width}}{B_{height}}$
We then compute the weight as
$w_{dist}=e^{\frac{-depth}{\lambda_{s}}}$
where $\lambda_{s}$ is a depth scaling factor. The predicted confidence score of each box is multiplied with $w_{dist}$, yielding higher confidence for boxes that appear closer in images. The smoothing module then uses the modified confidence score and finds the average modified confidence for each class in the track, along with their occurrence, selecting the class with the highest combination. Any track that is composed entirely of unknowns is dropped.
@@ III-C Open-Vocabulary Tracking Splits
Our method is evaluated in outdoor environments using custom data splits based on common object types and situations. Fig. 3 shows various statistics from nuScenes [22] and subsequent splits. One split follows the common open-vocabulary approach, treating rarer classes as novel. Another split differentiates between highway and urban objects based on the greater activity and mobility of some classes within urban areas. A final split focuses on objects with high variability in shape and size, represented by the average LiDAR points per object. These splits provide a robust framework for evaluating open-vocabulary tracking in realistic and diverse outdoor settings.
@@ IV Experiments
IV-A Datasets
Evaluation Metrics
IV-B Implementation Details
IV-C Benchmark Results
Table I shows results on the nuScenes [22] validation set using 3D proposals from CentrePoint [3] detector. We compare our open-vocabulary method on the various proposed splits with the baseline. For reference, we show the original closed-set results of 3DMOTFormer as an upper bound. The baseline method is a simple adaptation of 3DMOTFormer, where we replace class labels with labels obtained from 2D open-vocabulary detections and make use of confidence scores from the 2D detections as well. Baseline without unknowns drops any 3D proposals unmatched to 2D detections, while baseline with unknowns forms tracks with them and assigns them labels of the most common class within the track. For the first split, we show results for baseline with and without unmatched detections. With our presented adaptations, we achieve an overall AMOTA of 0.567 in the first split, with strong performance scores for the novel classes, improving significantly on top of the baseline. In each of the last two splits, we achieve a similar improvement over the baseline, with an overall AMOTA of 0.59 in split 2 and 0.536 in split 3. Particularly in split 3, we notice the performance being sub-par for novel class trucks; this is due to the vast variations in appearance, size, and type of vehicles for class trucks within the dataset, making this category a considerably difficult unknown class. We show some challenging novel class examples in Fig. 4.
@@ IV-C Benchmark Results
We evaluate our method’s generalization across different detectors, CenterPoint[3], MEGVII [25] and BEVFusion [26], as shown in Table II. Our approach delivers consistent performance across various 3D proposal sources. The second part of the table shows generalization results when trained on proposals from one detector and tested on another. The best results are from the standard detector, likely due to alignment between train and test bounding boxes. Interestingly, while BEVFusion gives the highest base class performance, it results in the lowest novel class performance, possibly due to focusing on high-quality base class detections and less noise for regulation effect.
@@ V Ablation
Class Agnostic GT Assignment: To build on top of our baseline we assign GT tracking IDs to 3D proposals using merely 3D IoU, omitting class labels. This improves results for the base classes at test time while results drop slightly for novel classes. This phenomenon occurs because class agnostic GT assignment improves target accuracy as compared to noisy open-vocabulary base class labels, leading to better learning on known classes, while reducing the regularization and dropping performance on unknown classes. However, the overall improvement in performance is significant.
@@ VI Conclusion
In conclusion, we introduce a novel approach for open-vocabulary 3D multi-object tracking, addressing the limitations of closed-set methods. By leveraging 2D open-vocabulary detections, and applying strategic adaptations to a 3D tracking framework such as class-agnostic tracking and confidence score prediction, we demonstrate improvements in tracking novel object classes. Our experiments highlight the robustness of our method. These contributions represent a meaningful step toward more flexible 3D tracking solutions for autonomous systems, enhancing their ability to generalize in dynamic environments.

### v2

@@ I INTRODUCTION
In this paper, we introduce the task of open-vocabulary 3D tracking, where trajectories of known and unknown object classes are estimated in the real 3D space by linking their positions across consecutive frames. We formulate the open-vocabulary tracking problem, propose splits to evaluate the performance across known and unknown classes, and introduce a method that bridges the performance gap between tracking known and unseen classes. Our approach leverages 2D open-vocabulary methods and 3D tracking frameworks to generalize to unseen categories effectively. Through extensive experiments, we demonstrate that our method achieves strong results through our adaptation strategies, providing a robust solution for open-vocabulary 3D tracking. Fig. 1 shows an example output of our solution. Our method can detect and track previously unseen object categories in three-dimensional space without relying on predefined labels.
@@ II Related Work
In 3D object tracking, various approaches have been explored, leveraging different features and methodologies. Traditional approaches rely on Kalman filters or other probabilistic models, with recent advances shifting towards deep learning-based methods. The tracking-by-detection paradigm remains dominant, particularly in 3D multi-object tracking (MOT). Model-based techniques, such as AB3DMOT [7], extend traditional Kalman filters to 3D for motion estimation, coupling it with the Hungarian method to provide a standard 3D MOT baseline. Another method, EagerMOT [8], combines observations from LiDAR and camera sensors, enhancing the ability to track distant objects and achieve precise trajectory localization. The Immortal Tracker [9] introduces a trajectory prediction mechanism using a Kalman filter to maintain tracks even when objects temporarily disappear from view, significantly reducing identity switches. CenterPoint [3] and CenterTrack [10] combine velocity estimates with point-based detection to track the centers of objects across multiple frames. Poly-MOT [1] introduces a method that adapts a tracking criterion based on object categories, employing category-specific motion models and a two-stage data association strategy. In 3DMOTFormer [2], the authors introduce a geometry-based 3D multi-object tracking framework built on the transformer architecture, using an Edge-Augmented Graph Transformer to perform data association through edge classification. While great advancement has been made in the 3D tracking task, all these methods rely on tracking a known predefined set of object classes.
@@ II-B Open-Vocabulary 3D
Open vocabulary 3D tasks are emerging in computer vision, driven by the need to generalize beyond predefined categories, particularly in dynamic environments. Recent advances in 2D open vocabulary tasks [11, 12, 13, 14] have inspired the extension of these concepts to the 3D domain, where the complexity of object shapes, occlusions, and sparse point-cloud data pose unique challenges. Introducing open-vocabulary techniques in 3D vision enables flexible, scalable models that identify diverse objects without extensive labeling. PointCLIP [15] leverages the Contrastive Vision-Language Pre-training (CLIP) model to achieve open-vocabulary point cloud recognition by aligning CLIP-encoded point clouds with 3D category texts through knowledge transfer from 2D images to 3D. For 3D open-vocabulary tasks, using image features obtained from vision language models and aligning them with point cloud features is common [16], [17]. OpenScene [18] generates dense features for 3D points embedded within the CLIP feature space by using image-per-pixel features from 2D semantic segmentation models [13]. OpenMask3D [19] introduces a zero-shot approach for open-vocabulary 3D instance segmentation by utilizing class-agnostic 3D instance masks and aggregating per-mask features through multi-view fusion of CLIP-based image embeddings. Beyond semantic segmentation, in [20], the authors propose combining a point cloud detector for object localization with cross-modal contrastive learning to connect an image, point cloud, and text representations, enabling open-vocabulary 3D detection in indoor scenes. Open-YOLO 3D [21] has also explored combining class-agnostic 3D proposals with strong 2D open-vocabulary detectors, demonstrating improvements in both accuracy and speed. However, open-vocabulary 3D tracking remains largely unexplored, presenting a significant gap in leveraging these emerging techniques for dynamic, real-time object tracking in the physical space.
@@ III-A Problem Formulation
Given a sequence of LiDAR frames $\mathbf{P}=\{P_{t}\}_{t=1}^{T}$, where $P_{t}$ represents the point cloud at time $t$, the goal of 3D multi-object tracking is to detect and track multiple objects $\{O_{i}^{t}\}_{i=1}^{N_{t}}$, where $O_{i}^{t}$ denotes the $i$-th object and $N_{t}$ are the total objects in frame $t$. Each $O_{i}^{t}$ comprises the object class $c$ and the 3D bounding box $b=[x,y,z,w,l,h,\theta]$ where $x,y,z$ represent the coordinates of the center of the bounding box, and $w,l,h$ are the width, length, and height of the bounding box, and $\theta$ is the heading angle, respectively.
@@ III-B Open-Vocabulary 3D Multi-Object Tracking
While 3DMOTFormer performs well for the closed set, it struggles to track unseen object classes due to its reliance on class-specific information during training. The tracker expects identified 3D boxes to generate edge proposals and then process those through the graph transformer. In a situation where it may receive a new label or an unlabelled object, the model will fail to track it. 3DMOTFormer also uses class-specific statistics within its pipeline, such as maximal class velocities, to compute association edges between tracks and new detections. Such information regarding unseen objects, that the system may encounter for the first time, is not available in a real-world open-vocabulary setting.
Class-Agnostic Tracking
To enable tracking irrespective of object class label, we modify the tracker to utilize only object position, dimensions, heading angle, and velocity as the initial features. Object detections from a given detector are treated as proposals (class-agnostic) where the class labels and confidence scores are dropped and not utilized.
To compute potential association edges, we replace class-specific distance thresholds with a fixed threshold, achieving a balance between maintaining true edges and avoiding unnecessary associations. We change the GT track ID assignment by computing 3D IoU between ground truth boxes $A^{train}$ and 3D detections belonging to $\mathcal{C}^{base}$ in a class-agnostic manner.
@@
As discussed earlier, we drop class labels and confidence scores from the detection features to ensure the tracker is independent of any information specific to object classes. As a result, the initial output from the system suffers from a lack of representation of the objectness score of each box. To overcome this issue, we use the detector objectness scores for $\mathcal{C}^{base}$ and train a regression head at the end of the tracker. The regression head is composed of a feed-forward network that takes as input processed detection features $h^{(Ld)}_{D,j}$ from the graph transformer and predicts a single confidence score. We employ a straightforward Mean Squared Error loss between predicted and true confidence scores to learn the scores for the detection boxes. The combined loss is $\mathcal{L}=\mathcal{L}_{m}+\lambda_{c}\mathcal{L}_{c}$, where $\mathcal{L}_{c}$ is the loss for confidence score, weighted by $\lambda_{c}$.
@@
While we train the tracker by removing all class-specific dependencies, we require a strong open-vocabulary classification for all object tracks, known and unknown, at test time. We leverage vision-language models to match regions in the images with text prompts, enabling it to detect and classify objects for which it has not been explicitly trained. 2D detections on multi-view images from the vehicle are obtained from an open-vocabulary 2D detector, which is prompted with all possible labels—both $\mathcal{C}^{base}$ and $\mathcal{C}^{novel}$. To enhance the cross-modal alignment between visual and textual information, we employ a YOLO-based framework, where a CLIP Encoder transforms input text into text embeddings while a darknet backbone extracts multi-scale image features. These representations are then refined using the RepVL-PAN module, which combines image and text embeddings to enable region-text matching.
The system allocates class labels to the 3D proposals by first projecting the 3D boxes onto multi-view images. The labels are then derived from the highest IoU overlap with 2D detections on the image plane. Since multi-view images provide a 360-degree view, a 3D box may project onto multiple images. A single projection is used, selecting its largest rectangular image region for IoU matching to obtain its 2D equivalent. Projections of the 3D detections that do not overlap with a 2D open-vocabulary detection are labeled ’unknown’. We explicitly identify these boxes at the end of the tracking process.
@@
At the output of the tracker, we apply a track smoothing module that addresses the issue of unknown boxes that do not match any 2D detections from the images, as well as improve label consistency throughout tracks to account for inaccuracies in the open-vocabulary detector outputs and errors accrued during projection from 3D to 2D. To calculate the most accurate class for each track, we first compute the weight of an object based on its detected bounding box in an image, with objects at a greater distance assigned a smaller weight. The depth is derived from the size of the bounding box, the object’s vertical position in the image for perspective correction, and the aspect ratio of the bounding box to ensure uniformity for all classes. Size is the normalized box area, given by: $B_{size}=\frac{B_{width}\times B_{height}}{I_{width}\times I_{height}}$
@@
where, $B$ refers to 2D box, $I$ refers to image, $y_{center}$ points to the vertical center of the box, $\alpha_{p}$ is the perspective correction factor estimated by comparing true depth to calculated depth, while $\beta_{ar}$ is the aspect ratio threshold selected by analyzing the distribution across common objects in outdoor scenes, and $aspect\_ratio=\frac{B_{width}}{B_{height}}$.
We then compute the weight as $w_{dist}=e^{\frac{-depth}{\lambda_{s}}}$ where $\lambda_{s}$ is a depth scaling factor. The predicted confidence score of each box is multiplied with $w_{dist}$, yielding higher confidence for boxes that appear closer in images. The smoothing module then uses the modified confidence score and finds the average modified confidence for each class in the track, along with their occurrence, selecting the class with the highest combination. Any track that is composed entirely of unknowns is dropped.
@@ III-C Open-Vocabulary Tracking Splits
Our method is evaluated in outdoor scenes using custom data splits based on common object types and situations. Fig. 3 shows various statistics from nuScenes [22] and subsequent splits. One split follows the common open-vocabulary approach, treating rarer classes as novel. Another split differentiates between highway and urban objects based on the greater activity of some classes within urban areas. A final split focuses on objects with high variability , represented by the average LiDAR points per object. These splits provide a robust framework for evaluating open-vocabulary tracking in realistic and diverse outdoor settings.
@@ IV Experiments
IV-A Datasets
Evaluation Metrics
IV-B Implementation Details
@@
IV-C Benchmark Results
Table I shows results on the nuScenes [22] validation set using 3D proposals from CentrePoint [3] detector. We compare our open-vocabulary method on the various proposed splits with the baseline. For reference, we show the original closed-set results of 3DMOTFormer as an upper bound. The baseline method is a simple adaptation of 3DMOTFormer, where we replace class labels with labels obtained from 2D open-vocabulary detections and make use of confidence scores from the 2D detections as well. For the first split, we show results for baseline with and without unmatched detections. With our presented adaptations, we achieve an overall AMOTA of 0.567 in the first split, with strong performance scores for the novel classes, improving significantly on top of the baseline. In each of the last two splits, we achieve a similar improvement over the baseline, with an overall AMOTA of 0.59 in split 2 and 0.536 in split 3. Particularly in split 3, we observe AMOTP increasing across some classes. This may be due to the Track Consistency Scoring, where unknown 3D proposals might be incorrectly identified, increasing location error but improving AMOTA due to better track continuity. We show some challenging novel class examples in Fig. 4.
We evaluate our method’s generalization across different detectors, CenterPoint[3], MEGVII [25] and BEVFusion [26], as shown in Table II. Our approach delivers consistent performance across various 3D proposal sources. The best results are from the standard detector, likely due to alignment between train and test bounding boxes. Interestingly, BEVFusion gives the highest base class performance and the lowest novel class performance, possibly due to training on high-quality base class proposals and less noise for regulation.
@@ V Ablation
Class-Agnostic GT Assignment: To build on top of our baseline we assign GT tracking IDs to 3D proposals using 3D IoU, omitting class labels. This improves results for the base classes at test time, while results drop slightly for novel classes. This occurs because class-agnostic GT assignment improves target accuracy as compared to noisy open-vocabulary base class labels, leading to better learning on known classes while reducing the regularization and dropping performance on unknown classes. However, the overall improvement in performance is significant.
@@ VI Conclusion
In conclusion, we introduce a novel approach for open-vocabulary 3D multi-object tracking, addressing closed-set limitations. By leveraging 2D open-vocabulary detections and applying strategic adaptations to a 3D tracking framework, such as class-agnostic tracking and confidence score prediction, we improve tracking novel object classes. Our experiments highlight the robustness of our method. These contributions represent a meaningful step toward more flexible 3D tracking solutions for autonomous systems, enhancing their ability to generalize in dynamic environments.
Acknowledgment The computations were enabled by resources provided by NAISS at Alvis partially funded by Swedish Research Council through grant agreement no. 2022-06725, LUMI hosted by CSC (Finland) and LUMI consortium, and by Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC.